OpenStudy (anonymous):

solve the equation,show your work 16^3=4^x

OpenStudy (math&ing001):

Have you seen the logarithmic function Ln yet ?

OpenStudy (jdoe0001):

\(\Large{ \bf {\color{blue}{ 16\implies 4^2\qquad \qquad 16^3\implies (4^2)^3}}\\ \quad \\ 16^3=4^x\implies (\square^\square)^\square =4^x}\)